# Review-Based Cross-Domain Recommendation via Hyperbolic Embedding and Hierarchy-Aware Domain Disentanglement

### v1

@@ 5.5. Scale Alignment and Disentanglement (RQ2)
In Figure 4, we describe the domain classification accuracy of the domain discriminator based on the application of scale alignment. Here, we employ three models: SER (Choi et al., 2022) which is a state-of-the-art disentanglement algorithm, HEAD (without Scale Alignment), and HEAD (with Scale Alignment). Since the domain label is binary (0 for source and 1 for target), we take the binary cross-entropy and set them as the y-axis. The x-axis is the training epochs. Both figures use the Luxury Beauty as the target domain, but each of them employs Clothing (left) and CDs (right) as the source domains, respectively. We discover that the discrimination accuracy is quite low in the left figure, where the two domains might be similar to each other. In addition to this, both figures represent that the discrimination accuracy of SER is better than HEAD (w/o SA), where HEAD has a larger scale than SER. This is because the hierarchical alignment in Eq. 14 has a separation characteristic. However, HEAD (w/ SA) achieves the best discrimination accuracy, which confirms the proposition 4.3.
@@ 5.6. Hierarchy Visualization (RQ3)
In Figure 5, we visualize item vectors in Digital Music dataset to assess the effect of hierarchy-aware embedding (Proposition 4.1). Here, we randomly sample 1,000 items and classify them based on their degrees. Specifically, we average the two item vectors $S_{i}$ and $\widehat{S}_{i}$ in Eq. 11 and project them onto the Poincaré ball (Eq. 1). The left figure employs simple root alignment (Eq. 13), while the right one further benefits from our degree-based normalization (Eq. 14). As observed in the left figure, nodes are quite randomly distributed regardless of their degrees. Although some nodes with higher degrees (red, $d>20$) are placed near the origin, nodes with lower degrees (purple, green, and blue) are positioned quite randomly. In contrast, the right figure shows that nodes are aligned based on degrees. From this, we conclude that the degree-based normalization successfully preserves the structural information, which leads to a better utilization of hyperbolic space eventually. For case studies, we highly recommend reading this article (Cao et al., 2022).
Figure 6. (RQ4) We describe the NDCG@10 of Video Games by varying the parameters $\lambda_{1}$ (x-axis, degree normalization) and $\lambda_{2}$ (y-axis, scale alignment) in Eq. 27, respectively
5.7. Ablation Study (RQ4)
Given the model with Poincaré Glove $HEAD^{*}_{\mathbb{P}}$, we vary the weights of the loss function in Eq. 27. Typically, the weight of the prediction loss (Eq. 26) is set to 1 since it is the main object of recommender systems. Now, we adjust the two hyper-parameters $\lambda_{1}$ and $\lambda_{2}$ that control the weight of hierarchy embedding and scale alignment. The experimental results are shown in Figure 6, where we conduct a grid search and plot the NDCG@10 score of the (Toys and Games, Video Games) pair using a heatmap. Here, the rows and columns represent $\lambda_{1}$ and $\lambda_{2}$, respectively. Firstly, we observe that the performances are dismal when the hyper-parameters take large values. This is because the ranking loss is overwhelmed by other functions, making the convergence of the parameters challenging. Instead, assigning small values for both $\lambda_{1}$ and $\lambda_{2}$ enhances the overall quality of recommendation by improving structural alignment and domain disentanglement. One might argue that the search for optimal parameters may require significant computational costs, but we find that suppressing these values below a specific threshold grants marginal improvements only. Please refer to (Choi et al., 2022) for the experimental results of using either domain-specific or shareable features.
@@ 6. CONCLUSION
Recent studies have addressed the challenge of data sparsity in recommender systems by integrating Cross-Domain Recommendation (CDR) with review texts. However, existing methods relying on an Euclidean space encounter difficulties due to the exponentially growing interactions between users and items. In response to this, we introduce a hyperbolic CDR as a potential solution and overcome several associated issues. Firstly, we identify some drawbacks related to root alignment and distance-based alignment, which are problematic in preserving the tree-like structure within a hyperbolic space, leading to a hierarchical collapse. To address these issues, we propose a novel solution: hierarchy-preserving embedding and domain disentanglement. Lastly, we provide a mathematical foundation to emphasize the theoretical relevance of our proposed strategies. Experimental results demonstrate the superiority of our model over state-of-the-art single and cross-domain algorithms.

### v3

@@ 5.3. Scale Alignment and Disentanglement (RQ2)
In Figure 3, we describe the domain classification accuracy of the domain discriminator based on the application of scale alignment. Here, we employ three models: SER (Choi et al., 2022) which is a state-of-the-art disentanglement algorithm, HEAD (without Scale Alignment), and HEAD (with Scale Alignment). Since the domain label is binary (0 for source and 1 for target), we describe the binary cross-entropy on the y-axis. The x-axis is the training epochs. Both figures use the Luxury Beauty as the target domain, but each of them employs Clothing (left) and CDs (right) as the source domains, respectively. We discover that the discrimination accuracy is quite low in the left figure, where the two domains are similar. In addition to this, both figures represent that the discrimination accuracy of SER is better than HEAD (w/o SA), where HEAD has a larger scale than SER. This is because the hierarchical alignment in Eq. 14 has a separation characteristic. However, HEAD (w/ SA) achieves the best discrimination accuracy, which confirms the proposition 4.3.
@@ 5.4. Hierarchy Visualization (RQ3)
In Figure 4, we visualize item vectors in Digital Music dataset to assess the effect of hierarchy-aware embedding (Proposition 4.1). Here, we randomly sample 1,000 items and classify them based on their degrees. Specifically, we average the two item vectors $S_{i}$ and $\widehat{S}_{i}$ in Eq. 11 and project them onto the Poincaré ball (Eq. 1). The left figure employs simple root alignment (Eq. 13), while the right one further benefits from our degree-based normalization (Eq. 14). As observed in the left figure, nodes are quite randomly distributed regardless of their degrees. Although some nodes with higher degrees (red, $d>20$) are placed near the origin, nodes with lower degrees (purple, green, and blue) are positioned quite randomly. In contrast, the right figure shows that nodes are aligned based on degrees. From this, we conclude that the degree-based normalization successfully preserves the structural information, which leads to a better utilization of hyperbolic space eventually. For case studies, we highly recommend reading this article (Cao et al., 2022a).
Figure 5. (RQ4) We describe the NDCG@10 of two datasets by varying the parameters $\lambda_{1}$ (x-axis, degree normalization) and $\lambda_{2}$ (y-axis, scale alignment) in Eq. 27, respectively
5.5. Parameter Sensitivity Analysis (RQ4)
Given the model with Poincaré Glove ${}^{\mathbb{H}}_{\mathbb{P}}$, we vary the weights of the loss function in Eq. 27. Typically, the weight of the prediction loss (Eq. 26) is set to 1 since it is the main object of recommender systems. Now, we adjust the two hyper-parameters $\lambda_{1}$ and $\lambda_{2}$ that control the weight of hierarchy embedding and scale alignment. The experimental results are shown in Figure 5, where we conduct a grid search and plot the NDCG@10 score through heatmap using the pairs of (CDs and Vinyl, Digital Music) and (Toys and Games, Video Games). Here, the rows and columns represent $\lambda_{1}$ and $\lambda_{2}$, respectively. Firstly, we observe that the performances are dismal when the hyper-parameters take large values. This is because the ranking loss is overwhelmed by other functions, making the convergence of the parameters challenging. Instead, assigning small values for both $\lambda_{1}$ and $\lambda_{2}$ enhances the overall quality of recommendation by improving structural alignment and domain disentanglement. As illustrated, we can see that setting $\lambda_{1}=\lambda_{2}=0.05$ in (a) Digital Music and $\lambda_{1}=0.1,\lambda_{2}=0.05$ in (b) Video Games achieve the best performance. One might argue that the search for optimal parameters may require huge computational costs, but we find that suppressing these values below a specific threshold grants marginal improvements only. Please refer to (Choi et al., 2022) for the ablation study of using either domain-specific or shareable features.
@@ 6. CONCLUSION
Recent studies have addressed the challenge of data sparsity in recommender systems by integrating Cross-Domain Recommendation (CDR) with review texts. However, existing methods relying on an Euclidean space encounter difficulties due to the exponentially growing interactions between users and items. In response to this, we introduce a hyperbolic CDR as a potential solution and overcome several associated issues. Firstly, we identify some drawbacks related to root- and distance-based alignment, which are problematic in preserving the tree-like structure within a hyperbolic space. To address these issues, we propose a novel solution: hierarchy-preserving embedding and domain disentanglement. Lastly, we provide a mathematical foundation to emphasize the theoretical relevance of our proposed strategies. Experimental results demonstrate the superiority of our model over state-of-the-art single and cross-domain algorithms.
7. Acknowledgments
This work was supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korean government (MSIT) (No. 2021-0-02068, IITP-2024-00156287), and KENTECH Research Grant (202200019A) funded by the National Research Foundation of Korea (NRF) (4199990214639).